Blossom Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $50 throughout the country to loyal alumni of over 2,100 schools. Blossom’s variable costs are 43% of sales; fixed costs are $114,000 per month.

Assume that variable costs increase to 45% of the current sales price and fixed costs increase by $12,000 per month. If Blossom were to raise its sales price 10% to cover these new costs, but the number of blankets sold were to drop by 5%, what would be the new annual operating income? (Round sales price to 2 decimal places, e.g. 52.75 and final answer to 0 decimal places, e.g. 5,275.)
Answer #1

Calculate new annual operating income

Total Per unit
Sales 109725 50*1.1 = 55
Variable cost 62194

21.5*1.45 = 31.175

Contribution margin 47531 23.825
Fixed cost 114000+12000 = 126000
Net operating income -78469